OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (2, -4) with a slope of 1/2 y = 1/2x + 5 y = 1/2x - 3 y = 1/2x - 5 y = 1/2x + 3

OpenStudy (anonymous):

@SpartanRace

OpenStudy (anonymous):

@Princer_Jones @princeharryyy

OpenStudy (anonymous):

(y-y1)=m(x-x1) is the equation of the line with slope m and the point (x1,y1) lies on it.

OpenStudy (anonymous):

C. 1/2-5

OpenStudy (anonymous):

1/2x-5 i mean

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

The tables below show the values of y corresponding to different values of x: Table A x 3 2 1 y 1 0 0 Table B x 3 5 4 y -2 1 1 Which statement is true for the tables? A. Both Table A and Table B represent functions. B. Both Table A and Table B do not represent functions. C. Table A does not represent a function, but Table B represents a function. D. Table A represents a function, but Table B does not represent a function.

OpenStudy (anonymous):

help anybody

OpenStudy (anonymous):

plz

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@Princer_Jones @princeharryyy

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

A.

OpenStudy (anonymous):

It's ok thank you

OpenStudy (michele_laino):

@IndianBoyNiko I think A, namely they are both functions

OpenStudy (princeharryyy):

yes u r right I got my concept wrong a bit @Princer_Jones

OpenStudy (princeharryyy):

I accept my mistake.... @Princer_Jones dats why I deleted what I wrote ....

OpenStudy (princeharryyy):

d answer is A

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

The graph of a function f(x) is shown below:

OpenStudy (anonymous):

What is the domain of f(x)? -2 < x ≤ 3 -2 ≤ x < 3 -1 < y ≤ 3 -1 ≤ y < 3

OpenStudy (anonymous):

-2 ≤ x < 3

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

The graph below shows four straight lines, P, Q, R, and S:

OpenStudy (anonymous):

Which line is represented by the function f(x) = 4x - 3? Line P Line Q Line R Line S

OpenStudy (anonymous):

@SpartanRace @Princer_Jones @EmmaMink @DanJS @KamiBug

OpenStudy (anonymous):

@campbell_st

OpenStudy (danjs):

hi, what is the question

OpenStudy (anonymous):

this one is tough bud.

OpenStudy (danjs):

Line S

OpenStudy (emmamink):

It would actually be y=1/2x-4, but since that is not an answer, then C would have to work.

OpenStudy (anonymous):

So which one is it, C or D

OpenStudy (michele_laino):

I think is the line S, because at x=0, y=-3, as required from your question

OpenStudy (anonymous):

The graph below shows the velocity f(t) of a runner during a certain time interval:

OpenStudy (anonymous):

Which of the following describes the intercepts on the graph? The initial velocity of the runner was 8 m/s, and the runner stopped after 2 seconds. The initial velocity of the runner was 2 m/s, and the runner stopped after 8 seconds. The initial acceleration of the runner was 8 m/s2, and the runner stopped after 2 seconds. The initial acceleration of the runner was 2 m/s2, and the runner stopped after 8 seconds.

OpenStudy (anonymous):

@EmmaMink

OpenStudy (anonymous):

@MathHomeworkHelp

OpenStudy (anonymous):

@Algorithmic

OpenStudy (anonymous):

@dayani.laforest

OpenStudy (anonymous):

@Samus_Aran

OpenStudy (anonymous):

@campbell_st

OpenStudy (danjs):

The initial velocity of the runner was 2 m/s and the runner stopped after 8 seconds.